OpenStudy (anonymous):

Consider the function with the formula g(x)=2x-11 a. what are the solution to the equation g(x) =-5 b.simplify each of the following g(x-2) and g(x)-2

OpenStudy (jmark):

Given g(x)=2x-11 a. what are the solution to the equation g(x) =-5 -5=2x-11 2x=-5+11 2x=6 x=3 b.simplify each of the following g(x-2) and g(x)-2 g(x-2)=2(x-2)-11 =2x-4-11 =2x-15 g(x)-2=2x-11-2 2x-9 CBSE Sample Papers for Class 9: http://cbse.edurite.com/cbse-sample-papers/cbse-sample-papers-for-class-9.html